OpenStudy (jennychan12):

How would you do integral from e to e^2 of 1/(x ln(x^8)) dx?

OpenStudy (jennychan12):

\[\int\limits_{e}^{e^2} \frac{ 1 }{ x \ln(x^8) } dx\]

zepdrix (zepdrix):

\[\large \int\limits_e^{e^2}\frac{1}{x \ln x^8}dx \qquad = \qquad \int\limits_e^{e^2}\frac{1}{8xln x}dx\]Remember your rules of logarithms? :) This will help quite a bit

OpenStudy (jennychan12):

oh wait. i forgot something

OpenStudy (abb0t):

\[\frac{ d }{ dx } \ln(x) =\]

OpenStudy (jennychan12):

waitttttt. i was stupid. i thought tjat e was a variable instead of the 2.7... number...

zepdrix (zepdrix):

Oh that's too funny :) lol

OpenStudy (jennychan12):

-_- that's just sad. :( i feel stupid. >:[

OpenStudy (abb0t):

Well all makes mistakes..... except me.

OpenStudy (jennychan12):

everybody makes mistakes :) -HM